The church youth group is planning a trip to Washington. A total of 127 people need rides. The church has available buses that hold 40 passengers, and several parents have volunteered station wagons that hold 8 passengers or sedans that hold 5 passengers. The youth leader is planning to use nine vehicles to transport the people. One parent said that if they didn't use any buses, tripled the number of station wagons, and doubled the number of sedans, they would be able to transport 126 people. How many buses, station wagons and sedans are they planning to use if they use 9 vehicles?
Where:
B = Bus
W = Wagon
S = Sedan
Indicate the algebraic equations used to solve this problem and then solve for the unknowns.
: The church youth group is planning a trip to Washington. A total of 127 people need rides. The church has available buses that hold 40 passengers, and several parents have volunteered station wagons that hold 8 passengers or sedans that hold 5 passengers. The youth leader is planning to use nine vehicles to transport the people. One parent said that if they didn't use any buses, tripled the number of station wagons, and doubled the number of sedans, they would be able to transport 126 people. How many buses, station wagons and sedans are they planning to use if they use 9 vehicles?
Where:
B = Bus
W = Wagon
S = Sedan
Indicate the algebraic equations used to solve this problem and then solve for the unknowns.

Answer by stanbon(18792)

The church youth group is planning a trip to Washington.
A total of 127 people need rides.
The church has available buses that hold 40 passengers,
and several parents have volunteered station wagons that hold 8 passengers
or sedans that hold 5 passengers.
The youth leader is planning to use nine vehicles to transport the people.
One parent said that if they didn't use any buses, tripled the number of station wagons, and doubled the number of sedans, they would be able to transport 126 people.
How many buses, station wagons and sedans are they planning to use if they use 9 vehicles?
Where:
B = # of Busses
W = # of Wagons
S = # of Sedans
Indicate the algebraic equations used to solve this problem and then solve for the unknowns.
--------------------
Vehicles: S + W + B = 9
People : 5S + 8W + 40B = 127
People : 2S + 3W + 0 = 126
---------------------------------------
I used the Matrix function of a TI calculator to get:
S = -81
W = 96
B = -5.9
---------
Comment: These answers dont make any sense. I suggest
you check the statement of your problem.
===================
Cheers,
stan H.